# AutoElicit: Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v4

@@ A.14 Empirically, is the language model approximating Bayesian inference?
A.15 AutoElicit with other language models
@@
A.15.1 Other OpenAI models
Figure 17: Test accuracy of the posterior distribution for some OpenAI language models. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@ A.15.1 Other OpenAI models
Figure 17 shows that GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in terms of their posterior performance in all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides prior distributions that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost. Further, in all tasks except Diabetes progression prediction, DeepSeek-R1-32B-Q4 provides better accuracy than GPT-4o-mini.
A.15.2 Open-source language models
Figure 18: Test accuracy of the posterior distribution for open-source language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@ A.15.2 Open-source language models
In contrast to the OpenAI models in Appendix A.15.1, these models are publicly available and open-source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, GPT-4-turbo and DeepSeek-R1-32B-Q4. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.7.3.
Figure 18 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of models. The 3B and 8B variants are significantly smaller in capacity to those in the main paper; justifying their inconsistent elicited priors. However, in some of the datasets they still produce good posterior performance.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with posterior performance similar to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo and comparable values to GPT-4 turbo or DeepSeek-R1-32B-Q4. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo and equal to DeepSeek-R1-32B-Q4. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for predictive tasks.
@@
A.16 Choosing between priors
As seen in Appendices A.15.1 and A.15.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that DeepSeek-R1-32B-Q4, GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task.

### v5

@@ A.15 Empirically, Is the Language Model Approximating Bayesian Inference?
A.16 AutoElicit with Other Language Models
@@
A.16.1 Other OpenAI Models
Figure 20: Test accuracy of the posterior distribution compared to GPT-4o mini. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on a log scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@ A.16.1 Other OpenAI Models
Figure 20 shows that GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in their posterior performance on all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides priors that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks, they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset, GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost. Further, in all tasks except Diabetes progression prediction, DeepSeek-R1-32B-Q4 provides better accuracy than GPT-4o-mini.
A.16.2 Open-source Language Models
Figure 21: Test accuracy of the posterior distribution for open-source language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on a log scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
@@ A.16.2 Open-source Language Models
In contrast to the OpenAI models in Appendix A.16.1, these models are publicly available and open-source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on-device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, GPT-4-turbo and DeepSeek-R1-32B-Q4. When sampling the posterior with the elicited priors, we use the same methods and options as given in Appendix A.7.3.
Figure 21 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of models. The 3B and 8B variants are significantly smaller in capacity to those in the main paper, justifying their inconsistent elicited priors. However, in some of the datasets, they still produce good posterior performance.
@@
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with posterior performance similar to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo and comparable values to GPT-4 turbo or DeepSeek-R1-32B-Q4. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo and equal to DeepSeek-R1-32B-Q4. Additionally, when predicting Breast Cancer, priors from Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the training sizes tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for predictive tasks.
@@
A.17 Choosing Between Priors
As seen in Appendices A.16.1 and A.16.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that DeepSeek-R1-32B-Q4, GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task.